# Dyna-Think: Synergizing Reasoning, Acting, and World Model Simulation in AI Agents

### v1

@@ Abstract
Recent progress in reasoning with large language models (LLMs), such as DeepSeek-R1, demonstrates impressive capabilities in domains like mathematics and coding, by exhibiting complex cognitive behaviors such as verification, goal decomposition, and self-reflection. However, it is unclear what behavior is effective and what behavior is missing for agentic tasks. In this work, we propose Dyna-Think, a thinking framework that integrates planning with an internal world model with reasoning and acting to enhance AI agent performance. To enable Dyna-Think, we propose Dyna-Think Imitation Learning (DIT) and Dyna-Think Dyna Training (DDT). To initialize a policy with Dyna-Think, DIT reconstructs the thinking process of R1 to focus on performing world model simulation relevant to the proposed (and planned) action, and trains the policy using this reconstructed data. To enhance Dyna-Think, DDT uses a two-stage training process to first improve the agent’s world modeling ability for next state prediction and critique generation, and then improve the agent’s action via policy training. We evaluate our methods on OSWorld, and demonstrate that Dyna-Think improves the agent’s in-domain and out-of-domain performance, achieving similar best-of-n performance compared to R1 while generating 2x less tokens on average. Our extensive empirical studies reveal that 1) using critique generation for world model training is effective to improving policy performance; and 2) improving a world model is effective to improving the performance of its AI agent. Our results suggest a promising research direction to integrate world-model simulation into AI agents to enhance their reasoning, planning, and acting capabilities.
@@ 1 Introduction
Autonomous AI agents powered by large language models (LLMs) have offered substantial promise in real-world applications, automating digital tasks such as software engineering (Jimenez et al., 2024; Wang et al., 2024b; Yang et al., 2024a), web navigation (Liu et al., 2018; Yao et al., 2023a; Zhou et al., 2024b; Koh et al., 2024a), computer-use (Xie et al., 2024; Anthropic, 2025b; Qin et al., 2025), and mobile device control (Rawles et al., 2023, 2024; Trivedi et al., 2024). As many computer use tasks require agents to interact with complex environments to achieve long-term objectives, one critical challenge is to reason and act efficiently over a large decision space.
Recent methods of test-time scaling (Snell et al., 2024) offer a potentially promising solution. For example, Yu et al. (2023, 2024b); Zhou et al. (2024a) show that LLM agents can significantly improve performance using search algorithms such as Monte Carlo Tree Search (MCTS) to allow additional interactions with the environment before decision-making. However, these methods require a large amount of expensive, time-consuming interactions with a real (or a separately learned) world model, limiting their applicability in real-world scenarios. Alternatively, many recent work (OpenAI, 2024; Valmeekam et al., 2024; Zhou et al., 2025) finds that LLM agents can also improve their performance by “thinking” longer, effectively internalizing parts of search into their reasoning process. For example, models such as OpenAI o1/o3 (OpenAI, 2024, 2025a), Claude-3.7-Sonnet (Anthropic, 2025a), and DeepSeek-R1 (DeepSeek-AI et al., 2025) generate significantly longer reasoning chains when faced with challenging tasks, exhibiting behaviors such as self-reflection, goal decomposition, verification, exploration, and more (Gandhi et al., 2025). Although many of these behaviors are crucial for solving challenging long-horizon tasks, recent work also finds that these models suffer from problems such as overthinking and fact-ignoring (Cuadron et al., 2025; Zhou et al., 2025), and are less efficient in certain tasks compared to non-reasoning models such as GPT-4o (Zhou et al., 2025). It is thus unclear what type of “thinking” is crucial to an agent’s performance for long-horizon tasks, and how to improve it via learning from experiences.
@@
The successes in Go, games, and robotics reveal that modeling and interacting with the environment is key to solving challenging long-horizon tasks. The Dyna algorithms (Sutton, 1991) offer a promising approach to combine real-world interactions with planning via simulation to improve an agent’s policy. However, modeling the whole environment for tasks such as computer-use (e.g., predicting screen content after entering ‘amazon.com’ in chrome) is challenging (Pathak et al., 2017; Wang et al., 2024a; Fang et al., 2025; Gu et al., 2025). In contrast, research in cognitive sciences (Marr, 1982; Rao & Ballard, 1999) shows that the human brain encodes a compressed representation of the external world, capturing only statistical regularities and meaningful structures related to current tasks. We thus propose Dyna-Think, a thinking framework that performs “compressed” world model simulation based on model-generated states and critiques (rewards), and integrates it with reasoning and acting to improve an AI agent’s performance. Figure 1 illustrates the difference between the Dyna-Think agent and DeepSeek-R1 using an example.
@@
The Dyna-Think framework consists of Dyna-Think Imitation Learning (DIT) and Dyna-Think Dyna Training (DDT). DIT improves a baseline agent policy by first distilling simulated experiences from R1-generated thinking tokens, and then revising the baseline policy using the simulated experiences via supervised learning. We finds models after DIT achieve a similar performance to R1-distilled models, but with 2X less thinking tokens on average. DDT is an extension to the Dyna-Q method (Sutton1998) that combines policy learning and world model training. Similar to Dyna-Q, DDT first constructs policy and world model training data using online rollouts in a real (digital) environment. Different from Dyna-Q, DDT is applicable to the problems with arbitrarily large state-action space by leveraging the prior knowledge encoded in a pre-trained LLM, and performs both policy learning and world model training based on a single LLM. During world model training, the model encodes state transitions by optimizing training objectives of next-state prediction and critiquing the states generated by its own. During policy learning, the agent’s actions are improved using successful rollouts via reinforcement learning.
@@
We evaluate Dyna-Think on the challenging OSWorld benchmark (Xie et al., 2024). Results show that Dyna-Think improves the agent’s in-domain and out-of-domain performance compared to only performing policy training (e.g., reinforcement learning) or training a separate world model (e.g., as in Dyna-Q), and that Dyna-Think leads to highly capable and cost-effective agents. For example, our 32B-parameter Dyna-Think model achieves a similar best-of-n performance to the much larger 685B-parameter R1 model, with 2x less tokens on average. A detailed analysis also reveals that 1) the critique-style world model training is effective in improving policy; and 2) the AI agents equipped with a stronger world model always perform better. Thus, our work demonstrates the potential of integrating planning and learning to develop future agents powered by reasoning models.
@@ 2 Related Work
Computer-use agents powered by large (multimodal) language models aim to automate tasks by controlling a computer, typically via GUI interactions with a virtual Ubuntu/Windows (Xie et al., 2024; Bonatti et al., 2024; Anthropic, 2025b). Early computer-use methods include reactive agents (Yao et al., 2023b; Xie et al., 2024) that directly prompts an LLM (e.g., GPT-4o) to make decisions on immediate observations without simulation or planning approaches. Recent advances include: 1) search-based methods (Zhou et al., 2024a; Koh et al., 2024b; Yu et al., 2024b) that augments LLMs with look-ahead search algorithms such as MCTS; and 2) hierarchical-planning methods that orchestrates multiple modules, tools, and LLMs to complete a task (Agashe et al., 2024, 2025; Liu et al., 2025; Gou et al., 2025; Yang et al., 2024b). However, search-based methods significantly increase inference time due to using additional interactions with the external environment; and hierarchical-planning methods often require complex human-designed rules and heuristics to coordinate multiple modules. We introduce our Dyna-Think framework to augment the thinking process of a single LLM agent by performing action-centric world model simulation.
@@ 3 Dyna-Think Framework
Figure 2: Our Dyna-Think framework integrates planning with world models simulation inside a single policy $\pi_{\mathcal{W}}(\theta)$, and performs both world model training and policy training with $\pi_{\mathcal{W}}(\theta)$.
We propose methods to 1) synergize reasoning, acting, and world modeling simulation in an agent’s thinking process; and 2) improve its policy and world modeling ability via Dyna-style training.
@@ 3.2 Dyna-Think Imitation Learning
Many expert LLMs capable of extensive thinking (e.g., DeepSeek R1, Claude-3.7-Sonnet, and OpenAI’s o3) exhibits complex behaviors during its thinking process. This includes being able to perform verification, self-reflection, error recovery, world modeling, and more. However, it is unclear which behaviors are critical for decision-making in long-horizon AI agent tasks. In our preliminary study, we find training on long CoT data with simulations/knowledge unrelated to the final action degrades performance, a phenomenon also found in other domains (Zhou et al., 2024c; Wu et al., 2025). To efficiently enable Dyna-Think, we propose Dyna-Think Imitation Learning (DIT), which reconstructs R1’s thinking process to only contain text related to reasoning, the final action, and the world modeling simulation related to the final action; and then trains a policy using the reconstructed data. To perform this reconstruction, we few-shot prompt GPT-4o (see Appendix C for examples and more details). After training, we refer to these models as $\pi_{\mathcal{W}}(\theta)$, highlighting their ability to perform world-model simulations during thinking.
@@ 3.3 Dyna-Think Dyna Training
Figure 3: Three different forms of world model data experimented in Dyna-Think training.
Despite DIT learning, $\pi_{\mathcal{W}}(\theta)$ at inference time can still make mistakes when facing unseen states unseen from training. However, creating a DIT training set that covers all possible states $|\mathcal{S}|$ is intractable. We propose Dyna-Think Dyna Training (DDT), a Dyna-style training method to perform policy and world model learning on $\pi_{\mathcal{W}}(\theta)$. Similar to Dyna, we first collect policy and world model learning data by performing $\pi_{\mathcal{W}}(\theta)$ rollouts in the real environment. Different from Dyna, we then directly perform both policy and world model learning on the same $\pi_{\mathcal{W}}(\theta)$ model. We illustrate our method in Figure 2.
@@
Given a rollout trajectory $\tau$, we also construct world model training dataset to train $\pi_{\mathcal{W}}(\theta)$ to model a function of the environment transition $\hat{\mathcal{T}_{f}}(o_{i},a_{i})$. We experiment with three different functions for $\hat{\mathcal{T}_{f}}$ in this work. $\hat{\mathcal{T}}(o_{i},a_{i})\to o_{i+1}$ which directly models the next state; $\hat{\mathcal{T}}_{\Delta}(o_{i},a_{i})\to\Delta(o_{i},o_{i+1}|a_{i})$ which trains to predict relevant changes in the next state caused by $a_{i}$; and $\hat{\mathcal{T}}_{\textrm{critic}}(o_{i},a_{i})\to\textrm{critic}(o_{i},a_{i}|o_{i+1})$ which trains to predict a critique generated by an LLM (GPT-4o) by comparing the world model simulations in $a_{i}$ with the previous and next state. An example world model simulation could be “After opening the terminal with Ctrl+Alt+T, type ‘cp dir2/hi.txt dir1/’ to copy …”, and an example critique is “Wait, maybe we need to first ensure that ‘dir1’ exists in the current directory…”. We illustrate these three functions in Figure 3. For more details, please refer to Section D.1.
Following Dyna methods, DDT training performs two-stage learning by first training $\pi_{\mathcal{W}}(\theta)$ on the world model dataset, and then training $\pi_{\mathcal{W}}(\theta)$ on the policy dataset. For other implementation details such as prompts used by GPT-4o to generate a feedback and how feedback data is formatted during training, please refer to Appendix D.
@@ 4.1 Benchmarks
We evaluate our methods on OSWorld (Xie et al., 2024), a diverse benchmark consisting of 369 open-ended computer tasks that involves real web and desktop apps in open domains, OS file I/O, and workflow tasks. Tasks are categorized into 10 different domains based on different desktop applications involved, such as OS Terminal, LibreOffice Calc, LibreOffice Impress, LibreOffice Writer, Chrome, VLC Player, Thunderbird, VS Code, GIMP and Workflow.
To evaluate the self-improving ability of AI agents, many methods (Huang et al., 2022; Yu et al., 2024a; Chen et al., 2025; Jin et al., 2025) consider domains where the model has non-trivial initial performance. However, in OSWorld we find many domains such as LibreOffice Calc and Workflow to be extremely challenging, with current models achieving less than 10% success rate or solving only 2-3 tasks often due to state representation issues (see Section E.1). To this end, we evaluate our method on 5 domains that are more accessible to existing models, including OS, Chrome, VS Code, GIMP, and Thunderbird.

### v3

@@ Abstract
Recent progress in reasoning with large language models (LLMs), such as DeepSeek-R1, demonstrates impressive capabilities in domains like mathematics and coding, by exhibiting complex cognitive behaviors such as verification, goal decomposition, and self-reflection. However, it is unclear what behavior is effective and what behavior is missing for agentic tasks. In this work, we propose Dyna-Think, a thinking framework that integrates planning with an internal world model with reasoning and acting to enhance AI agent performance. To enable Dyna-Think, we propose Dyna-Think Imitation Learning (DIT) and Dyna-Think Dyna Training (DDT). To initialize a policy with Dyna-Think, DIT reconstructs the thinking process of R1 to focus on performing world model simulation relevant to the proposed (and planned) action, and trains the policy using this reconstructed data. To enhance Dyna-Think, DDT uses a two-stage training process to first improve the agent’s world modeling ability for next state prediction and critique generation, and then improve the agent’s action via policy training. We evaluate our methods on OSWorld and WindowsAgentArena, and demonstrate that Dyna-Think improves the agent’s in-domain and out-of-domain performance, achieving similar best-of-n performance compared to R1 while generating 2x less tokens on average. Our extensive empirical studies reveal that 1) using critique generation for world model training is effective to improving policy performance; and 2) improving a world model is effective to improving the performance of its AI agent. Our results suggest a promising research direction to integrate world-model simulation into AI agents to enhance their reasoning, planning, and acting capabilities.
@@ 1 Introduction
Autonomous AI agents powered by large language models (LLMs) have offered substantial promise in real-world applications, automating digital tasks such as software engineering (Jimenez et al., 2024; Wang et al., 2024b; Yang et al., 2024a), web navigation (Liu et al., 2018; Yao et al., 2023a; Zhou et al., 2024b; Koh et al., 2024a), computer-use (Xie et al., 2024; Anthropic, 2025b; Qin et al., 2025), and mobile device control (Rawles et al., 2023; 2024; Trivedi et al., 2024). As many computer use tasks require agents to interact with complex environments to achieve long-term objectives, one critical challenge is to reason and act efficiently over a large decision space.
Recent methods of test-time scaling (Snell et al., 2024) offer a potentially promising solution. For example, Yu et al. (2023; 2024b); Zhou et al. (2024a) show that LLM agents can significantly improve performance using search algorithms such as Monte Carlo Tree Search (MCTS) to allow additional interactions with the environment before decision-making. However, these methods require a large amount of expensive, time-consuming interactions with a real (or a separately learned) world model, limiting their applicability in real-world scenarios. Alternatively, many recent work (OpenAI, 2024; Valmeekam et al., 2024; Zhou et al., 2025) finds that LLM agents can also improve their performance by “thinking” longer, effectively internalizing parts of search into their reasoning process. For example, models such as OpenAI o1/o3 (OpenAI, 2024; 2025a), Claude-3.7-Sonnet (Anthropic, 2025a), and DeepSeek-R1 (DeepSeek-AI et al., 2025) generate significantly longer reasoning chains when faced with challenging tasks, exhibiting behaviors such as self-reflection, goal decomposition, verification, exploration, and more (Gandhi et al., 2025). Although many of these behaviors are crucial for solving challenging long-horizon tasks, recent work also finds that these models suffer from problems such as overthinking and fact-ignoring (Cuadron et al., 2025; Zhou et al., 2025), and are less efficient in certain tasks compared to non-reasoning models such as GPT-4o (Zhou et al., 2025). It is thus unclear what type of “thinking” is crucial to an agent’s performance for long-horizon tasks, and how to improve it via learning from experiences.
@@
The successes in Go, games, and robotics reveal that modeling and interacting with the environment is key to solving challenging long-horizon tasks. The Dyna algorithms (Sutton, 1991) offer a promising approach to combine real-world interactions with planning via simulation to improve an agent’s policy. However, modeling the whole environment for tasks such as computer-use (e.g., predicting screen content after entering ‘amazon.com’ in chrome) is challenging (Pathak et al., 2017; Wang et al., 2024a; Fang et al., 2025; Gu et al., 2025). In contrast, research in cognitive sciences (Marr, 1982; Rao & Ballard, 1999) shows that the human brain encodes a compressed representation of the external world, capturing only statistical regularities and meaningful structures related to current tasks. We thus propose Dyna-Think, a thinking framework that performs “compressed” world model simulation based on model-generated states and critiques (rewards), and integrates it with reasoning and acting to improve an AI agent’s performance. Figure˜1 illustrates the difference between the Dyna-Think agent and DeepSeek-R1 using an example.
@@
The Dyna-Think framework consists of Dyna-Think Imitation Learning (DIT) and Dyna-Think Dyna Training (DDT). DIT improves a baseline agent policy by first distilling simulated experiences from R1-generated thinking tokens, and then revising the baseline policy using the simulated experiences via supervised learning. We finds models after DIT achieve a similar performance to R1-distilled models, but with 2X less thinking tokens on average. To further improve DIT, we propose DDT, is an extension to the Dyna-Q method (Sutton & Barto, 2018) that combines policy and world model learning during online training. Similar to Dyna-Q, DDT first constructs policy and world model training data using online rollouts in a real (digital) environment. Different from Dyna-Q, DDT is applicable to the problems with arbitrarily large state-action space by leveraging the prior knowledge encoded in a pre-trained LLM, and performs both policy learning and world model training based on a single LLM. During world model training, the model encodes state transitions by optimizing training objectives of next-state prediction and critiquing the states generated by its own. During policy learning, the agent’s actions are improved using successful rollouts via reinforcement learning.
@@
We evaluate Dyna-Think on OSWorld (Xie et al., 2024) and WindowsAgentArena (Bonatti et al., 2024). Results show that Dyna-Think improves the agent’s in-domain and out-of-domain performance compared to only performing policy training (e.g., reinforcement learning) or training a separate world model (e.g., as in Dyna-Q), and that Dyna-Think leads to highly capable and cost-effective agents. For example, our 32B-parameter Dyna-Think model achieves a similar best-of-n performance compared to the much larger 685B-parameter R1 model, with 2x less tokens on average. Detailed analysis also reveals that 1) the critique-style world model training is effective in improving policy; and 2) AI agents with a stronger world model achieves better performance. Thus, our work demonstrates the potential of integrating planning and learning to develop future agents powered by reasoning models.
@@ 2 Related Work
Computer-use agents powered by large (multimodal) language models aim to automate tasks by controlling a computer, typically via GUI interactions with a virtual Ubuntu/Windows (Xie et al., 2024; Bonatti et al., 2024; Anthropic, 2025b). Early computer-use methods include reactive agents (Yao et al., 2023b; Xie et al., 2024) that directly prompts an LLM (e.g., GPT-4o) to make decisions on immediate observations without simulation or planning approaches. Recent advances include: 1) search-based methods (Zhou et al., 2024a; Koh et al., 2024b; Yu et al., 2024b) that augments LLMs with look-ahead search algorithms such as MCTS; and 2) hierarchical-planning methods that orchestrates multiple modules, tools, and LLMs to complete a task (Agashe et al., 2024; 2025; Liu et al., 2025; Gou et al., 2025; Yang et al., 2024b). However, search-based methods significantly increase inference time due to using additional interactions with the external environment; and hierarchical-planning methods often require complex human-designed rules and heuristics to coordinate multiple modules. We introduce our Dyna-Think framework to augment the thinking process of a single LLM agent by performing action-centric world model simulation.
@@ 3 Dyna-Think Framework
Figure 2: Our Dyna-Think framework synergizes planning with world model simulation in an agent’s reasoning process, and performs both world model training and policy training with $\pi_{\mathcal{W}}(\theta)$.
Accurate world simulation is challenging for AI agents. We propose Dyna-Think, a two-stage training method to address this. First, we introduce DIT (Section˜3.2) to synergize reasoning, acting, and world model simulation in an agent’s thinking process via imitation learning. Then, we introduce DDT (Section˜3.3) to further improve its policy and world modeling ability via Dyna-style training.
@@ 3.2 Dyna-Think Imitation Learning
Many expert LLMs capable of extensive thinking (e.g., DeepSeek R1, Claude-3.7-Sonnet, and OpenAI’s o3) exhibits complex behaviors during its thinking process. This includes being able to perform verification, self-reflection, error recovery, world modeling, and more. However, it is unclear which behaviors are critical for decision-making in long-horizon AI agent tasks. In our preliminary study, we find training on long CoT data with simulations/knowledge unrelated to the final action degrades performance, a phenomenon also found in other domains (Zhou et al., 2024c; Wu et al., 2025). To enable Dyna-Think for “weak” non-reasoning models, we propose Dyna-Think Imitation Learning (DIT) to first construct reasoning data that emphasizes on world simulation and perform imitation learning. Specifically, DIT reconstructs R1’s thinking process to only contain text related to reasoning, the final action, and the world modeling simulation related to the final action; and then trains a policy using the reconstructed data. To perform this reconstruction, we few-shot prompt GPT-4o (see Appendix˜D for examples and more details). After training, we refer to these models as $\pi_{\mathcal{W}}(\theta)$, highlighting their ability to perform world-model simulations during thinking.
@@ 3.3 Dyna-Think Dyna Training
Despite DIT learning, $\pi_{\mathcal{W}}(\theta)$ at inference time can still make mistakes when facing unseen states unseen from training. However, creating a DIT training set that covers all possible states $|\mathcal{S}|$ is intractable. To further improve DIT-trained models, we propose Dyna-Think Dyna Training (DDT), a Dyna-style method that performs both policy and world model learning during online training. Similar to Dyna, we first collect policy and world model learning data by performing $\pi_{\mathcal{W}}(\theta)$ rollouts in the real environment. Different from Dyna, we then directly perform both policy and world model learning on the same $\pi_{\mathcal{W}}(\theta)$ model. We illustrate our method in Figure˜2.
@@
Figure 3: Three different forms of world model data experimented in DDT.
@@
Given a rollout trajectory $\tau$, we also construct world model training dataset to train $\pi_{\mathcal{W}}(\theta)$ to model a function of the environment transition $\hat{\mathcal{T}_{f}}(o_{i},a_{i})$. We experiment with three different functions for $\hat{\mathcal{T}_{f}}$ in this work. $\hat{\mathcal{T}}(o_{i},a_{i})\to o_{i+1}$ which directly models the next state; $\hat{\mathcal{T}}_{\Delta}(o_{i},a_{i})\to\Delta(o_{i},o_{i+1}|a_{i})$ which trains to predict relevant changes in the next state caused by $a_{i}$; and $\hat{\mathcal{T}}_{\textrm{critic}}(o_{i},a_{i})\to\textrm{critic}(o_{i},a_{i}|o_{i+1})$ which trains to predict a critique generated by an LLM (GPT-4o) by comparing the world model simulations in $a_{i}$ with the previous and next state. An example world model simulation could be “After opening the terminal with Ctrl+Alt+T, type ‘cp dir2/hi.txt dir1/’ to copy …”, and an example critique is “Wait, maybe we need to first ensure that ‘dir1’ exists in the current directory…”. We illustrate these three functions in Figure˜3. To utilize these objectives to enhance the world modeling ability of the policy model, we optimize next-state prediction $\hat{\mathcal{T}}$ and state-changes prediction $\hat{\mathcal{T}}_{\Delta}$ as an “auxiliary” task trained by standard language modeling loss alongside policy training; and we optimize $\hat{\mathcal{T}}_{\textrm{critic}}$ by injecting the generated critique back into the action ($a_{i}^{\prime}=a_{i}\oplus\textrm{critic}$) and then train the policy to predict critic tokens in $a_{i}^{\prime}$ using language modeling loss. Intuitively, $\hat{\mathcal{T}}$ and $\hat{\mathcal{T}}_{\Delta}$ enhance a policy’s world modeling ability implicitly, while $\hat{\mathcal{T}}_{\textrm{critic}}$ is more explicit. For more details, please see Sections˜E.1 and E.2.
Finally, we combine policy and world model training, DDT follows Dyna methods and perform two-stage learning by first training $\pi_{\mathcal{W}}(\theta)$ on the world model dataset, and then training $\pi_{\mathcal{W}}(\theta)$ on the policy dataset. For pseudocode, please see Algorithm˜1. For other implementation details such as critique prompts and how training data is formatted, please refer to Appendix˜E.
@@ 4.1 Benchmarks
We mainly evaluate our methods on OSWorld (Xie et al., 2024), a diverse benchmark consisting of 369 open-ended computer tasks that involves real web and desktop apps in open domains, OS file I/O, and workflow tasks. Tasks are categorized into 10 different domains based on different desktop applications involved, such as OS Terminal, LibreOffice Calc, LibreOffice Impress, LibreOffice Writer, Chrome, VLC Player, Thunderbird, VS Code, GIMP and Workflow.
To evaluate the self-improving ability of AI agents, many methods (Huang et al., 2022; Yu et al., 2024a; Chen et al., 2025; Jin et al., 2025) consider domains where the model has non-trivial initial performance. However, in OSWorld we find many domains such as LibreOffice Calc and Workflow to be extremely challenging, with current models achieving less than 10% success rate or solving only 2-3 tasks often due to state representation issues (see Section˜F.1). To this end, we evaluate our method on 5 domains that are more accessible to existing models, including OS, Chrome, VS Code, GIMP, and Thunderbird; and also additionally on WindowsAgentArena (Bonatti et al., 2024) to further measure model’s generalization ability to a different operating system (Windows OS).